OpenStudy (anonymous):

What is the volume of the box if it is scaled down by a factor of 1/10? Length: 9in. Width: 5in. Height: 3in.

OpenStudy (anonymous):

Length = 9/10 Width = 5/10 Height = 3/10 Volume = Length*Height*Width

OpenStudy (anonymous):

but idk what to do with those fractions?? @thushananth01

OpenStudy (anonymous):

Multiply the fractions which will result in the volume

OpenStudy (anonymous):

27/200 uhh

OpenStudy (anonymous):

@thushananth01

OpenStudy (anonymous):

Yes if your calculation is right